Suppose that there are only two goods produced in the economy, Call center services and banking services. Prices (P), quantities (Q) and the number of workers (W) occupied in the production of each good for year 1 and year 2 are given by the following table : Year 1 Year 2 Two products P1 Q1 W1 P2 Q2 W2 Call centre 10 100 50 12 100 50 Banking 10 200 50 12 230 60 a. What is nominal GDP in each year? 


1
Expert's answer
2021-11-30T17:42:29-0500

Solution:

a.). Nominal GDP is derived by multiplying the current year quantity output by the current market price.

Call center services:

Year 1 = 10 x 100 = 1,000

Banking Services:

Year 1 = 10 x 200 = 2,000

Nominal GDP Year 1 = 1,000 + 2,000 = 3,000

 

Call center services:

Year 2 = 12 x 100 = 1,200

Banking Services:

Year 2 = 12 x 230 = 2,760

Nominal GDP Year 1 = 1,200 + 2,760 = 3,960